Which fraction is found between 7/12 and 5/6 on a number line? A.)13/24 B.)11/12 C.)5/12 D.)17/24
Answer by Alan3354(69443)

Which fraction is found between 7/12 and 5/6 on a number line? A.)13/24 B.)11/12 C.)5/12 D.)17/24
------------
Hint: 7/12 = 14/24 & 5/6 = 20/24